# The Causal Effect of the Two-For-One Strategy in the National Basketball Association

### v1

@@ 1 Introduction
In recent years, the use of data and statistical methods to analyze sports strategies has gained significant traction, driven by the availability of detailed game data and advances in data science. In professional sports leagues like the National Basketball Association (NBA), decision-makers seek to optimize team performance through strategic adjustments informed by empirical analysis. Among these strategies, the “two-for-one” tactic, where a team attempts to gain an extra possession by timing shots near the end of a quarter, has become popular. However, despite its widespread use, little rigorous evidence exists on whether this strategy provides a measurable advantage. This study aims to address this gap by applying causal inference methods to determine the effectiveness of the two-for-one strategy.
An increasing body of research has applied causal inference methods to investigate the causal effects of various sports strategies. For example, Gibbs et al. (2022) utilize causal inference to evaluate the impact of calling a timeout to disrupt an opposing team’s scoring run in NBA games. Nakahara et al. (2023) analyze pitching strategies in baseball through a causal inference framework, while Yam & Lopez (2019) examine the decision-making process surrounding fourth-down plays in the NFL, aiming to quantify the effectiveness of these high-stakes decisions. Additionally, Wu et al. (2021) explore the effect of crossing the ball in soccer, seeking to determine whether this strategy significantly influences scoring opportunities.
@@
Causal inference (Imbens & Rubin, 2015) methods are essential tools for rigorously analyzing the effects of treatments or interventions in observational settings, particularly when randomized controlled trials (RCTs) are impractical, as is often the case in sports. In this study, we focus on estimating the average treatment effect (ATE) of the NBA’s two-for-one strategy. The ATE quantifies the average difference in outcomes (e.g., game score or quarter lead) between situations where the two-for-one strategy is employed and those where it is not. Estimating the ATE in observational data, however, requires addressing confounding variables—factors that influence both the treatment and the outcome, potentially biasing the results. To mitigate such bias, we apply inverse probability weighting and propensity score matching, two widely used methods in causal inference that allow for more reliable estimation of treatment effects.
Inverse probability weighting involves estimating the probability (propensity score) that a given observation will receive the treatment (the two-for-one strategy) based on observed covariates (e.g., player lineup, time in the game). Each observation is then weighted by the inverse of its propensity score to create a pseudo-population with a similar distribution of covariates between treated and control groups. Alternatively, matching pairs treated and control observations that share similar values of the propensity score, effectively matching the treatment and control groups on key covariates. This approach further reduces the risk of confounding and allows for a more direct comparison of the outcomes between the two groups, simulating the conditions of an RCT. Together, these methods help isolate the causal effect of the two-for-one strategy from other confounding influences.
@@
Beyond estimating the ATE, it is crucial to assess whether the treatment effect varies across different contexts or player configurations. To do this, we employ the causal forest (Athey et al., 2019; Wager & Athey, 2018) framework, a machine learning-based method that enables the estimation of heterogeneous treatment effects (HTE). Traditional ATE estimation assumes that the treatment effect is constant across all observations, but in real-world settings like the NBA, the effectiveness of a strategy is likely to vary based on game context, player lineups, and other factors. The causal forest framework addresses this by partitioning the data into subgroups using decision trees (Hastie et al., 2009). Each subgroup represents a distinct set of conditions where the treatment effect may differ. Causal forests construct many such trees and aggregate the resulting estimates to produce a flexible, non-parametric estimate of treatment effect heterogeneity. By identifying subgroups with distinct treatment effects, causal forests allow us to uncover whether the two-for-one strategy is more or less effective in different situations, such as when certain players are on the court or at different points in a game.
@@
In addition to estimating treatment effect heterogeneity, we apply the concept of rank average treatment effect (RATE) (Yadlowsky et al., 2024) to better understand the strategy’s impact. RATE works by ranking observations based on their predicted treatment effects, identifying which game contexts or player configurations are most likely to benefit from the two-for-one strategy. The RATE method quantifies the expected impact on the top ranked subgroups and compares these with the average effect across all observations. This ranking approach is particularly valuable in practice, as it highlights the specific conditions under which the two-for-one strategy is most advantageous, offering actionable insights for coaches and analysts looking to optimize in-game decisions.
1.1 Contributions
This study leverages detailed play-by-play data from the 2018-19 and 2021-22 NBA seasons, including variables such as player lineups, betting odds, and player ratings, to rigorously analyze the effectiveness of the two-for-one strategy. By framing the two-for-one strategy as a causal inference problem, we provide a formal statistical model that allows for reliable estimation of treatment effects while controlling for confounding factors. The methods used in this analysis—inverse probability weighting, matching, causal forests, and RATE—are all designed to ensure that our results are robust and generalizable to other contexts in sports analytics.
The main contributions of this paper are as follows:
  1. 1.
    We frame the two-for-one strategy as a causal inference problem, translating a complex sports decision into a statistical model that can be rigorously analyzed.
  2. 2.
    We estimate the average treatment effect of the two-for-one strategy, providing quantitative evidence of its effectiveness in NBA games.
  3. 3.
    We investigate potential heterogeneity in the strategy’s impact using the causal forest framework, finding that the strategy’s effectiveness appears relatively consistent across various game contexts.
Our findings contribute to sports analytics by offering data-driven insights into an influential NBA strategy, providing coaches and analysts with valuable, evidence-based guidance for optimizing in-game tactics.
@@ 2 Definitions and Data
The TFO (Feng, 2015; Fischer, 2015) strategy in the NBA is a time-management tactic that aims to maximize a team’s scoring opportunities by securing two offensive possessions at the end of a quarter, while limiting the opponent to just one. This strategy is commonly employed in the last 30–40 seconds of a quarter, especially during close games, where maximizing scoring chances can be crucial. The TFO strategy is primarily focused on managing the game clock to optimize possessions and increase scoring potential before the quarter ends.
To implement the TFO, a team aims to take a shot early enough in the shot clock so that even if the opponent uses a full possession, the team can secure the ball again before time expires. For example, if a team has possession with 36 seconds remaining in the quarter, they may attempt a quick, high-quality shot within 6–8 seconds. This leaves approximately 30 seconds on the clock, and even if the opposing team takes nearly the entire 24 seconds allowed per possession, the original team would regain possession with around 5–10 seconds remaining, allowing them one last shot before the quarter ends.
@@ 2.1 Definitions
In order to determine the causal effect of the TFO strategy in the NBA, we need a well-defined treatment. That is, we need to be able to classify scenarios during an NBA game as those in which a team enacted the TFO strategy, and those in which they did not. This is not a trivial matter. Teams do not announce that they are attempting a TFO, and it is unclear how attune players are to the game and shot clock during any particular play. Thus we rely on a combination of time and play outcomes in order to identify TFO opportunities (TFO-O). These are scenarios when a team comes into possession of the ball with enough time left in the quarter such that taking a shot or initiating offense relatively early in the shot clock leads to a increased chance to get an extra possession, while waiting until the end of the shot clock to do these things increases the chance the opposing team will have the last possession. We then categorize those TFO-O into attempts (TFO-A) and non-attempts (TFO-NA).
@@
Several important characteristics of NBA game play inform these definitions. The shot clock in the NBA is 24 seconds, meaning that from the time a team gains possession of the ball after a turnover, rebound, or made basket, the team has 24 seconds in which to attempt a shot. If a team gains possession with less than 24 seconds left in the quarter, the shot clock is turned off and the team in possession can attempt to have the last possession of the quarter by waiting until the very end to take a shot. Therefore, if a team wants to improve its chances of gaining the last possession of the quarter, it should hope that the opposing team gains possession with more than 24 seconds left.
@@
However, it can take up to a few seconds for possession to change, particularly after a missed shot. As an example, suppose a player on Team A takes a shot with 6 seconds left in the quarter. It takes one second for the ball to get from the player’s hand to the rim, another second to bounce on the rim a couple of times before coming to the ground, and a third second before a player on Team B possesses it, now with only 3 seconds left in the quarter. Furthermore, if there is too little time left in the quarter when a team gains possession, that team has a very small chance of scoring any points.
@@
With these characteristics in mind, we set a limit of 28 seconds left in the quarter, and say that if a team takes a shot after this point and the possession changes, it is unlikely that they will get a quality possession at the end of the quarter. Working backwards, then, a team needs to gain possession of the ball with enough time before the 28 seconds mark in order to potentially attempt a reasonable shot by this point. We set this lower limit to be 35 seconds. We choose 43 seconds as an upper limit on when a team can gain possession to exclude those scenarios where a team can get to this optimal shot time simply by waiting until the shot clock for their possession is close to 0. Thus we define a TFO-O as one in which a team comes into possession of the ball with between 35 and 43 seconds left in the quarter.
@@
If a team presented with a TFO-O takes a shot or is fouled while taking a shot with at least 28 seconds to go in the quarter, we consider that a clear TFO-A. Other play results are harder to categorize. We also consider an opportunity a TFO-A if the team is fouled with at least 28 seconds to go in the quarter, with the rationale being that fouls more often occur when a team is trying to initiate offense. On the other hand, we consider an opportunity to be a TFO-NA if the team maintains possession until there are less than 28 seconds left in the quarter without taking a shot or being fouled.
@@
We remove all other play outcomes that occur during a TFO-O from consideration rather than defining them as a TFO-A or a TFO-NA. The most common of these is a turnover. Turnovers can happen while a team is initiating offense, such as a charge foul, or just trying to hold the ball until the end of the shot clock, such as a steal that happens far away from the basket. Due to the lack of information on turnovers available in play-by-play data, we choose not to put them into either the TFO-A or TFO-NA categories. The same is true for other, less common plays such as a jump ball, a defensive three-second violation, and a kicked ball violation.
We define our response variable as the Post Opportunity Difference (POD), denoted by $Y_{i}$. The POD quantifies the effectiveness of the TFO strategy by measuring the difference in score differential from the time of the TFO-O to the end of the quarter. To calculate $Y_{i}$, we first determine the score difference between the teams at the time of the TFO-O, denoted $Y_{i,TFO-O}$. Next, we record the score difference at the end of the quarter, denoted $Y_{i,EOQ}$. The POD is then defined as:
$Y_{i}=Y_{i,EOQ}-Y_{i,TFO-O}$ (1)
@@
A positive value of $Y_{i}$ indicates that the team increased its score margin by the end of the quarter. Conversely, a negative $Y_{i}$ implies that the team’s score margin decreased after the TFO-O. Thus if the TFO strategy is effective the POD would be higher for teams that attempt a TFO than those who do not.
@@ 2.2 Data
The main source of data we use is the play-by-play data from the NBA. This is a list of all the different plays that took place in an NBA game as well as supplemental information such as the players involved in the play, the score of the game at the time of the play, the quarter the play took place in, and the amount of time left in that quarter. We obtain play-by-play data for each regular season game in the 2018-2019 and 2021-2022 seasons using the nbastatR package (Bresler, 2024). The 2018-2019 season was the last regular season before the Covid-19 pandemic altered the NBA schedule, and the 2021-2022 season was the first full season after the pandemic, making these natural seasons to consider.
Using this play-by-play data, we identify all TFO-O in the first three quarters of each NBA regular season game. We choose to exclude the fourth quarters because the behavior of NBA games is often much different at the end of those quarters. If the game is not close, teams are no longer necessarily worried about maximizing their possessions because of the feeling that nothing they do will affect the outcome. If the game is very close, teams tend to foul more to try to regain possession. We believe that excluding these scenarios allows us to better isolate the treatment. For the 2018-2019 season, there were a total of 1036 TFO-A and 529 TFO-NA. For the 2021-2022 season there were a total of 950 TFO-A and 462 TFO-NA.
@@
For player-specific information, we use individual player ratings from the NBA 2K video game series (HoopsHype, 2018, 2021), which provide relative strength assessments for each player. We obtain starting lineups from BasketballReference.com (Basketball Reference, n.d.) and use this information, combined with play-by-play data, to determine which players are on the court at any given time. These covariates allow us to account for variations in team and player strength as well as game tempo, enhancing the robustness of our analysis. Table 1 provides the name and definitions of the variables we use.
@@
Spread The betting point spread for the game.
Total Score The projected total score from betting odds.
@@
In the following section, we describe the causal inference methodology we employed, detailing our computation of the average treatment effect and our investigation into treatment effect heterogeneity.

### v3

@@ Abstract
Executive Summary
The two-for-one basketball strategy involves attempting to take a shot early in a possession in order to maximize the number of possessions at the end of a period. Based on our analysis of two seasons of data from the National Basketball Association, we estimate that teams that attempt a two-for-one strategy gain slightly more than half a point per attempt, on average, compared to teams that do not attempt a two-for-one strategy. We do not find that differences in certain relevant covariates lead to substantial differences in this point gain. Our analysis thus empirically validates the growing consensus of the effectiveness of the two-for-one strategy in the National Basketball Association.
@@ 1 Introduction
In recent years, the use of data and statistical methods to analyze sports strategies has gained significant traction, driven by the availability of detailed game data and advances in data science. In professional sports leagues like the National Basketball Association (NBA), decision-makers seek to optimize team performance through strategic adjustments informed by empirical analysis. Among these strategies, the “two-for-one” tactic, where a team attempts to gain an extra possession by timing shots near the end of a period, has become popular. However, despite its widespread use, little rigorous evidence exists on whether this strategy provides a measurable advantage. This study aims to address this gap by applying causal inference methods to determine the effectiveness of the two-for-one strategy.
An increasing body of research has applied causal inference methods to investigate the causal effects of various sports strategies. For example, Gibbs et al. (2022) utilize causal inference to evaluate the impact of calling a timeout to disrupt an opposing team’s scoring run in NBA games. Nakahara et al. (2023) analyze pitching strategies in baseball through a causal inference framework, while Yam and Lopez (2019) examine the decision-making process surrounding fourth-down plays in the NFL, aiming to quantify the effectiveness of these high-stakes decisions. Additionally, Wu et al. (2021) explore the effect of crossing the ball in soccer, seeking to determine whether this strategy significantly influences scoring opportunities.
@@
Causal inference (Imbens and Rubin, 2015) methods are essential tools for rigorously analyzing the effects of treatments or interventions in observational settings, particularly when randomized controlled trials (RCTs) are impractical, as is often the case in sports. In this study, we focus on estimating the average treatment effect (ATE) of the NBA’s two-for-one strategy. The ATE quantifies the average difference in outcomes (e.g., difference in score margin) between situations where the two-for-one strategy is employed and those where it is not. Estimating the ATE in observational data, however, requires addressing confounding variables—factors that influence both the treatment and the outcome, potentially biasing the results. To mitigate such bias, we apply augmented inverse probability weighting, a widely used method in causal inference that allows for more reliable estimation of treatment effects. This method involves estimating the probability (propensity score) that a given observation will receive the treatment (the two-for-one strategy) based on observed covariates (e.g., player lineup, time in the game). Each observation is then weighted by the inverse of its propensity score to create a pseudo-population with a similar distribution of covariates between treated and control groups.
@@
Beyond estimating the ATE, it is crucial to assess whether the treatment effect varies across different contexts or player configurations. To do this, we employ the causal forest (Athey and Wager, 2019; Wager and Athey, 2018) framework, a machine learning-based method that enables the estimation of heterogeneous treatment effects (HTE). Traditional ATE estimation assumes that the treatment effect is constant across all observations, but in real-world settings like the NBA, the effectiveness of a strategy is likely to vary based on game context, player lineups, and other factors. The causal forest framework addresses this by partitioning the data into subgroups using decision trees (Hastie et al., 2009). Each subgroup represents a distinct set of conditions where the treatment effect may differ. Causal forests construct many such trees and aggregate the resulting estimates to produce a flexible, non-parametric estimate of treatment effect heterogeneity. By identifying subgroups with distinct treatment effects, causal forests allow us to uncover whether the two-for-one strategy is more or less effective in different situations, such as when certain players are on the court or at different points in a game.
@@
In addition to estimating treatment effect heterogeneity, we apply the concept of rank average treatment effect (RATE) (Yadlowsky et al., 2025) to better understand the strategy’s impact. RATE works by ranking observations based on their predicted treatment effects, identifying which game contexts or player configurations are most likely to benefit from the two-for-one strategy. The RATE method quantifies the expected impact on the top ranked subgroups and compares these with the average effect across all observations. This ranking approach is particularly valuable in practice, as it highlights the specific conditions under which the two-for-one strategy is most advantageous, offering actionable insights for coaches and analysts looking to optimize in-game decisions.
@@ 2 Definitions and Data
The TFO (Feng, 2015; Fischer, 2015) strategy in the NBA is a time-management tactic that aims to maximize a team’s scoring opportunities by securing two offensive possessions at the end of a period, while limiting the opponent to just one. This strategy is commonly employed in the last 30–40 seconds of a period, especially during close games, where maximizing scoring chances can be crucial. The TFO strategy is primarily focused on managing the game clock to optimize possessions and increase scoring potential before the period ends. To implement the TFO, a team aims to take a shot with a certain amount of time remaining in the period so that even if the opponent uses a full possession, the team can secure the ball again before time expires.
In order to attempt a TFO, a team may be required to take a shot early in the possession, regardless of the quality of shot. This runs counter to the common strategy in basketball of progressing through an offensive set until an opportunity for a quality shot is realized. Thus the potential down side of attempting a TFO is that even if a team ends up with more possessions at the end of the period than their opponent, they may not ultimately benefit because of poorer scoring opportunities.
@@ 2.1 Definitions
To see if the potential gains of a TFO strategy outweigh the potential losses we need to determine a causal effect, and for that we need a well-defined treatment. That is, we need to be able to classify scenarios during an NBA game as those in which a team enacted the TFO strategy, and those in which they did not. This is not a trivial matter. Teams do not announce that they are attempting a TFO, and it is unclear how attune players are to the game and shot clock during any particular play. Thus we rely on a combination of time and play outcomes in order to identify TFO opportunities. These are scenarios when a team comes into possession of the ball with enough time left in the period such that taking a shot or initiating offense relatively early in the shot clock leads to a increased chance to get an extra possession, while waiting until the end of the shot clock to do these things increases the chance the opposing team will have the last possession. We then categorize those TFO opportunities into attempts (treatment group) and non-attempts (control group).
@@
Several important characteristics of NBA game play inform these definitions. NBA games are played in four 12-minute periods. Possession at the beginning of the game is determined by a jump ball, the result of which also determines which team gets possession at the start of subsequent periods. The shot clock in the NBA is 24 seconds, meaning that from the time a team gains possession of the ball after a turnover, defensive rebound, or made basket by the opposing team, the team has 24 seconds in which to attempt a shot. If a team rebounds their own missed shot the shot clock is reset to 14 seconds instead of 24. If a team gains possession with less than 24 seconds left in the period, the shot clock is turned off and the team in possession can attempt to have the last possession of the period by waiting until the very end to take a shot. Therefore, if a team wants to improve its chances of gaining the last possession of the period, it should hope that the opposing team gains possession with more than 24 seconds left.
@@
However, it can take up to a few seconds for possession to change, particularly after a missed shot. As an example, suppose a player on Team A takes a shot with 6 seconds left in the period. It takes one second for the ball to get from the player’s hand to the rim, another second to bounce on the rim a couple of times before coming to the ground, and a third second before a player on Team B possesses it, now with only 3 seconds left in the period. Furthermore, if there is too little time left in the period when a team gains possession, that team has a very small chance of scoring any points.
@@
With these characteristics in mind, we set a limit of 28 seconds left in the period, and say that if a team takes a shot after this point and the possession changes, it is unlikely that they will get a quality possession at the end of the period. Working backwards, then, a team needs to gain possession of the ball with enough time before the 28 seconds mark in order to potentially attempt a reasonable shot by this point. We set this lower limit to be 35 seconds. We choose 43 seconds as an upper limit on when a team can gain possession to exclude those scenarios where a team can get to this optimal shot time simply by waiting until the shot clock for their possession is close to 0. Thus we define a TFO opportunity as one in which a team comes into possession of the ball with between 35 and 43 seconds left in the period.
@@
If a team presented with a TFO opportunity takes a shot or is fouled while taking a shot with at least 28 seconds to go in the period, we consider that a clear TFO attempt. Other play results are harder to categorize. We also consider an opportunity an attempt if the team is fouled with at least 28 seconds to go in the period, with the rationale being that fouls more often occur when a team is trying to initiate offense. On the other hand, we consider an opportunity to be a non-attempt if the team maintains possession until there are less than 28 seconds left in the period without taking a shot or being fouled.
@@
We remove all other play outcomes that occur during a TFO opportunity from consideration rather than defining them as either an attempt or a non-attempt. The most common of these is a turnover. Turnovers can happen while a team is initiating offense, such as a charge foul, or just trying to hold the ball until the end of the shot clock, such as a steal that happens far away from the basket. Due to the lack of information on turnovers available in play-by-play data, we choose to remove them from our analysis. The same is true for other, less common plays such as a jump ball, a defensive three-second violation, and a kicked ball violation. Figure 1 provides a visual description of our treatment and control groups.
Figure 1: Diagram of the treatment definition.
We believe this to be a reasonable treatment definition for the TFO strategy. However, it also has limitations. One of these is that the timing cutoffs, while well motivated, are nevertheless somewhat arbitrary. Thus in Section 4.1.4 we present a sensitivity analysis in which we adjust these cutoffs slightly to explore the robustness of our estimates to changes in the definition.
A second limitation is that, as mentioned earlier, there is no way to know the intent of the coaches or the players on the floor with respect to the TFO strategy, which forces us to rely on event data for our treatment definition. There are certainly scenarios when a player intends to implement a TFO strategy but is not part of our treatment group because, for example, they lost control of the ball while going up for a shot and it is recorded as a turnover. The distinction from the clinical trials literature between intent-to-treat (ITT) and per-protocol (PP) analyses is relevant here. In the former case subjects are analyzed based on the original intended treatment, where in the latter they are analyzed based on the treatment they actually receive. Molero-Calafell et al. (2024) discuss the strengths and limitations of these analyses in relation to whether a study is more explanatory, meaning executed under more ideal experimental conditions, or pragmatic, executed under more real world conditions. Certainly our study is more pragmatic by their definitions, and thus they would recommend an ITT analysis with perhaps a PP analysis as a complement. However, in this framework we do not know the original intent and thus our approach is more similar to that of a PP analysis because an ITT analysis is not possible. However, Molero-Calafell et al. (2024) mention that employing advanced statistical methods such as the inverse probability weighting we use in our analysis can help enhance the validity of PP estimates.
@@
We define our response variable as the Post Opportunity Difference (POD), which measures the difference in score differential from the time of the TFO opportunity to the end of the period. A positive POD indicates that the team increased its score margin by the end of the period. Conversely, a negative POD implies that the team’s score margin decreased after the TFO opportunity. Thus if the TFO strategy is effective the POD would be higher for teams that attempt a TFO than those who do not.
@@ 2.2 Data
The main source of data we use is play-by-play data from the NBA. This is a list of all the different plays that took place in an NBA game as well as supplemental information such as the players involved in the play, the score of the game at the time of the play, the period the play took place in, and the amount of time left in that period. We obtain play-by-play data from the nbastatR package (Bresler, 2024), which interfaces with the NBA Stats API and other reference sources to give a variety of data and summary statistics at the player, team, game, and season level. This includes play-by-play data dating back to the 1996-1997 season, the first year the NBA kept such data. For our analysis we focus on two recent seasons, that of 2018-2019 and 2021-2022. The 2018-2019 season was the first regular season under a modified set of shot clock rules as well as the last regular season before the Covid-19 pandemic altered the NBA schedule, and the 2021-2022 season was the first full season after the pandemic, making these natural seasons to consider in order to explore the overall effect of the TFO strategy as well as possible heterogeneity across multiple seasons.
Using this play-by-play data, we identify all TFO opportunitys in the first three periods of each NBA regular season game. We choose to exclude the fourth periods and overtime because the behavior of NBA games is often much different at the end of those periods. If the game is not close, teams are no longer necessarily worried about maximizing their possessions because of the feeling that nothing they do will affect the outcome. If the game is very close, teams tend to utilize a strategy of intentionally fouling in order to increase the number of possessions in the game. We believe that excluding these scenarios allows us to better isolate the treatment, and we present a numerical justification for this choice in Section 4.1.5. We also exclude TFO opportunities in which the team has already had an earlier TFO opportunity in the same period. For the 2018-2019 season, there were a total of 1567 TFO opportunities, of which 1036 were classified as attempts and the remaining 529 classified as non-attempts. For the 2021-2022 season there were a total of 1479 TFO opportunities, of which 950 were classified as attempts and the remaining 462 classified as non-attempts. Each NBA season contains 1230 regular season games, so there was an average between 1 and 2 TFO opportunities per game based on our criteria.
@@
For player-specific information, we use individual player ratings from the NBA 2K video game series (HoopsHype, 2018, 2021) which provide relative strength assessments for each player. While not perfect, these ratings are a reasonable proxy for the overall ability of players. In addition to previous years’ statistics, ratings are informed by expert projections, particularly for players new to the league (Irving, 2021; Panerio, 2024). We obtain starting lineups from BasketballReference.com (Basketball Reference, ) and use this information, combined with play-by-play data, to determine which players are on the court at any given time. These covariates allow us to account for variations in team and player strength as well as game tempo, enhancing the robustness of our analysis. Table 1 provides the name and definitions of the variables we use.
@@
Spread The projected difference in final score for the two teams, set by oddsmakers.
Total Score The projected sum of the scores of the two teams, set by oddsmakers.
@@
Combining these covariates and the play-by-play data into a single usable data set requires a substantial amount of data cleaning. The play-by-play data includes substitutions but not a list of the current players on the floor, so we have to input the starting lineups and then create variables that track the current lineup on the floor. Player names are also represented differently in the various sources which complicates data merging. Finally, the play-by-play data provides narrative descriptions such as the type of shot, which is a challenge when categorizing plays. As an example, in Table 2 described in the section below, the phrases jump shot, finger roll layup, and dunk all indicate a shot has been attempted.
2.3 Example
To illustrate the definitions and data described above, consider an example from a game played between the Phoenix Suns and the Golden State Warriors on October 22, 2018. Table 2 shows several rows of the play-by-play data with a subset of relevant columns. In the first period of the game the Suns gained possession with 42 seconds left after a Warriors turnover, giving the Suns a TFO opportunity. However, they did not take a shot early in the shot clock. Rather, they committed a turnover with 25 seconds left. By our definition, this is a non-attempt. The Warriors gained possession with the turnover and made a 3 point shot, after which there was no more scoring in the period. So from the time of the TFO opportunity until the end of the period the Suns scored 0 points and the Warriors scored 3 points, for a POD of -3.
With 37 seconds left in the second period of the same game, the Warriors gained possession after a made shot by the Suns, giving the Warriors a TFO opportunity. With 31 seconds left Kevin Durant made a dunk for the Warriors, and we consider this shot attempt with more than 28 seconds to go in the period a TFO attempt. The Suns gained possession at that point and committed a turnover with 12 seconds left. Thus the Warriors gained a final possession, and Quinn Cook made a three-point shot with one second left, which did not allow the Suns to take another shot. From the time of the TFO opportunity until the end of the period, the Warriors scored 5 points and the Suns scored 0 points, for a POD of 5.
Pr. Time DescriptionPlayHome DescriptionPlayVisitor ScoreH ScoreV
1 0:42 Curry Out of Bounds - NA NA NA
Bad Pass Turnover
Turnover (P3.T5)
1 0:25 NA Booker Discontinue NA NA
Dribble Turnover (P2.T3)
1 0:25 SUB: Cook FOR Looney NA NA NA
1 0:25 SUB: Thompson FOR NA NA NA
McKinnie
1 0:07 Jerebko 27’ 3PT Step NA 32 23
Back Jump Shot (3 PTS)
(Thompson 2 AST)
1 0:03 Iguodala STEAL (1 STL) Booker Lost Ball Turnover NA NA
1 0:01 MISS Curry 55’ 3PT NA NA NA
Jump Shot
1 0:$0$ WARRIORS Rebound NA NA NA
2 0:37 NA Booker 4’ Driving Finger 65 47
Roll Layup (14 PTS)
2 0:31 Durant 3’ Driving Dunk NA 67 47
(15 PTS)
2 0:12 NA Booker Out of Bounds NA NA
Lost Ball Turnover (P5.T9)
2 0:12 SUB: Cook FOR Jones NA NA NA
2 0:12 NA SUB: Bridges FOR NA NA
Anderson
2 0:01 Cook 24’ 3PT Pullup NA 70 47
Jump Shot (4 PTS)
(Thompson 3 AST)
Table 2: Section of play-by-play data obtained from the nbastatR package. The variables shown are the period of play, the time left in that period, descriptions of the play for both the home and visitor teams, and score for both the home and visitor teams. Variable names have been altered.